### v1

# Novel View Synthesis from a Single Image with Pretrained Diffusion Guidance

Abstract.
Recent 3D novel view synthesis (NVS) methods are limited to single-object-centric scenes generated from new viewpoints and struggle with complex environments. They often require extensive 3D data for training, lacking generalization beyond training distribution. Conversely, 3D-free methods can generate text-controlled views of complex, in-the-wild scenes using a pretrained stable diffusion model without tedious fine-tuning, but lack camera control. In this paper, we introduce HawkI++, a method capable of generating camera-controlled viewpoints from a single input image. HawkI++ excels in handling complex and diverse scenes without additional 3D data or extensive training. It leverages widely available pretrained NVS models for weak guidance, integrating this knowledge into a 3D-free view synthesis approach to achieve the desired results efficiently. Our experimental results demonstrate that HawkI++ outperforms existing models in both qualitative and quantitative evaluations, providing high-fidelity and consistent novel view synthesis at desired camera angles across a wide variety of scenes.
3D-free, View Consistency, Background Inclusion, Guidance Models
journal: TOGccs: Computing methodologies Reconstructionccs: Computing methodologies Computational photography
Figure 2. Method. HawkI++ generates a high fidelity camera controlled novel viewpoint of a single image $I_{input}$, its text description and designated angle information. It infuses prior information from pre-trained NVS models into the text to image stable diffusion architecture in a 3D-free training and inference procedure.
1. Introduction
3D-free methods, such as HawkI (Kothandaraman et al., 2023b), can generate aerial views from complex, in-the-wild input images without needing additional multi-view or 3D information for fine-tuning or inference. Specifically, HawkI facilitates aerial view image synthesis by combining Inverse Perspective Mapping (IPM) (Kothandaraman et al., 2023a), a classical computer vision technique, with sampling methods inspired by information theory within the text-to-image stable diffusion backbone model. While it can handle any in-the-wild complex scene, it lacks the ability to achieve precise control over camera angles, when generating novel-view images.
Can we combine the benefits of both 3D-driven approaches and 3D-free methods to generate camera-controlled novel views at specified azimuth and elevation angles from a single input image? Additionally, can we achieve such using 3D-based pretrained novel view synthesis models alongside pretrained text-to-image diffusion models, without requiring 3D datasets or extensive training?
In this paper, we present HawkI++, a novel approach that generates camera-controlled novel views at specified azimuth and elevation angles from a single input image, without requiring 3D datasets or extensive training. Our method leverages information from pretrained 3D-based models like Zero123++ (Shi et al., 2023) alongside the pretrained stable diffusion model, following the 3D-free HawkI-style approach to achieve the desired goal. We provide extensive qualitative and quantitative comparisons to demonstrate effectiveness of our method.
Figure 3. Results on HawkI-Syn. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 4. Results on HawkI-Real. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ View-CLIP $\uparrow$ CLIPD $\uparrow$ View-CLIPD $\uparrow$ CLIP-I $\uparrow$
HawkI-Syn $(30^{\circ},30^{\circ})$ Ours 0.5667 29.9125 19.7100 0.0534 0.0568 0.8334
HawkI-Syn $(30^{\circ},30^{\circ})$ HawkI 0.5998 27.6516 18.6122 0.0405 0.0549 0.8108
HawkI-Syn $(30^{\circ},30^{\circ})$ Zero123++ 0.5793 29.8709 18.4740 0.0552 0.0540 0.8330
HawkI-Syn $(30^{\circ},30^{\circ})$ Stable Zero123 0.6853 20.5335 18.0746 0.0528 0.0520 0.6181
HawkI-Syn $(30^{\circ},270^{\circ})$ Ours 0.5791 29.2654 19.1836 0.0553 0.0611 0.8274
HawkI-Syn $(30^{\circ},270^{\circ})$ HawkI 0.5996 29.0022 19.1088 0.0321 0.0498 0.8438
HawkI-Syn $(30^{\circ},270^{\circ})$ Zero123++ 0.5911 28.9917 19.1491 0.0482 0.0597 0.7917
HawkI-Syn $(30^{\circ},270^{\circ})$ Stable Zero123 0.6441 21.2231 18.3738 0.0406 0.0403 0.6152
HawkI-Real $(30^{\circ},30^{\circ})$ Ours 0.6075 30.9333 20.7728 0.0829 0.0730 0.8333
HawkI-Real $(30^{\circ},30^{\circ})$ HawkI 0.6405 28.5431 18.5430 0.0506 0.0468 0.8302
HawkI-Real $(30^{\circ},30^{\circ})$ Zero123++ 0.6679 29.8661 19.9530 0.0624 0.0467 0.8049
HawkI-Real $(30^{\circ},30^{\circ})$ Stable Zero123 0.6624 22.8604 18.7536 0.0523 0.0658 0.6740
HawkI-Real $(30^{\circ},270^{\circ})$ Ours 0.6180 30.0677 19.6756 0.0652 0.0709 0.8350
HawkI-Real $(30^{\circ},270^{\circ})$ HawkI 0.6497 28.4452 18.4231 0.0574 0.0577 0.8236
HawkI-Real $(30^{\circ},270^{\circ})$ Zero123++ 0.6935 27.2773 19.5926 0.0605 0.0698 0.7212
HawkI-Real $(30^{\circ},270^{\circ})$ Stable Zero123 0.6466 23.6915 17.8817 0.0593 0.0658 0.6726
Table 1. Quantitative Results. Evaluation on six metrics demonstrate the superior results of our method over prior work.
2. HawkI++
We present HawkI++ (Figure 2), a model crafted to generate camera-controlled novel-view images from a single input image ($I_{input}$) and its text description ($t_{input}$) (e.g., “An ancient Egyptian pyramid in the desert,” obtained using the BLIP-2 model). HawkI++ eliminates the need for training data, 3D data, or multi-view data. Instead, it utilizes a pretrained text-to-2D image stable diffusion model as a strong prior, along with pretrained novel-view synthesis (NVS) models like Zero123++ for guidance. Our method combines information from the pretrained NVS model and performs a rapid inference-time optimization and inference routine, following a HawkI-style inference approach, to generate novel-view images of any given in-the-wild complex input scene at specified azimuth ($\alpha_{azi}$) and elevation angles ($\alpha_{elev}$).
Inference-time Optimization
Inspired by HawkI (Kothandaraman et al., 2023b) that utilizes Inverse Perspective Mappingas weak guidance for aerial-view generation (Kothandaraman et al., 2023a), we employ a pretrained NVS model $G$ to obtain a coarse prediction, $I_{view}$, of $I_{input}$ at ($\alpha_{azi}$, $\alpha_{elev}$). This prediction is represented as $I_{view}=G(I_{input},\alpha_{azi},\alpha_{elev})$. Although $I_{view}$ is not a fully accurate depiction of the desired target, it provides weak or pseudo guidance for the model regarding the content and direction of the desired viewpoint transformation. Subsequently, we utilize the pretrained text-to-image stable diffusion model to perform inference-time optimization using a HawkI-style inference.
Specifically, initially, we enhance the CLIP embedding for $t_{input}$ with $I_{input}$ to derive $e_{optim}$, which is optimized to most accurately reconstruct $I_{input}$:
[…]
Subsequently, the LoRA layers (with parameters $\theta_{LoRA}$) within the cross-attention layers of the diffusion model are fine-tuned at $e_{optim}$ to replicate $I_{input}$, employing the diffusion denoising loss function,
[…]
while keeping the rest of the UNet model unchanged. This process is then iterated for $I_{view}$, where $e_{optim}$ is again refined to best reconstruct $I_{view}$. Following this, the LoRA layers are adjusted to capture the nuances of the weak guidance image $I_{view}$, guiding the transformation towards the desired outcome.
Inference
To generate the camera-controlled image with designated azimuth ($\alpha_{azi}$) and elevation angles ($\alpha_{elev}$), we use the target text description $t_{target}$, which varies according to the corresponding $\alpha_{azi}$ and $\alpha_{elev}$. For instance, if $\alpha_{azi}=30^{\circ}$ and $\alpha_{elev}=30^{\circ}$, $t_{target}$ can be formatted as ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees” + $t_{input}$ (e.g., ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees, An ancient Egyptian pyramid in the desert.”). Next, we use the finetuned diffusion model to generate the target image using $t_{target}$, along with mutual information guidance (Kothandaraman et al., 2023b), which enforces similarity between the contents of the generated and input images.
3. Experiments and Results
Datasets
We utilize the HawkI-Syn (Kothandaraman et al., 2023b) and HawkI-Real (Kothandaraman et al., 2023b) datasets that feature complex scenes with multiple foreground objects and background.
Baselines
We compare our method with state-of-the-art view synthesis methods, including Zero123++ (Shi et al., 2023), HawkI (Kothandaraman et al., 2023b) and Stable Zero123.
Implementation Details
We employ the stable diffusion 2.1 model as the backbone for all our experiments and results. To generate the pseudo guidance images for different viewpoints, we use the pretrained Zero123++ (Shi et al., 2023) model. The hyperparameters for finetuning and inference are set in accordance with HawkI (Kothandaraman et al., 2023b).
3.1. Qualitative analysis
We evaluate our method on four distinct viewpoints:
$\{(\alpha_{azi},\alpha_{elev})\mid(30^{\circ},30^{\circ}),(-20^{\circ},210^{\circ}),(30^{\circ},270^{\circ}),\vspace*{-0.5em}(-20^{\circ},330^{\circ})\}.$
We present qualitative representative results and comparisons with Zero123++ (Shi et al., 2023), HawkI (Kothandaraman et al., 2023b), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 3 and 4. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 3 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, HawkI++ accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints.
Similar observations are made for HawkI-Real results shown in Figure 4. StableZero123 is ineffective. Zero123++ fails to capture background or detailed information. For example, when tasked with camera control for an image of the Eiffel Tower, Zero123++ focuses solely on the Eiffel Tower, ignoring surrounding details. The original HawkI model, while producing aerial views, fails in angle conversion tasks. In contrast, HawkI++ accurately performs angle conversion tasks at $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$, including the Seine River in the background for the Eiffel Tower image, showcasing its superiority. Camera control tasks for the HawkI-Real dataset, including images like the Hawaii beach and a cat, further demonstrate HawkI++’s excellence compared to other models. The key benefits of HawkI++ over 3D-based NVS methods such as Zero123++ and 3D-free methods such as HawkI arises by merging the strengths of 3D-based techniques into a 3D-free optimization process, effectively combining the best features of both.
3.2. Quantitative Evaluation
Following prior work (Shi et al., 2023; Kothandaraman et al., 2023b; Liu et al., 2023), we evaluate our method using six metrics - (i) LPIPS (Zhang et al., 2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score (Radford et al., 2021) and View-CLIP Score: Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{\text{azi}},\alpha_{\text{elev}})$ viewpoint, while the View-CLIP Score focuses specifically on the viewpoint. Higher values are preferred. (iii) CLIP-I (Ruiz et al., 2023), CLIPD, and View-CLIPD: Evaluate image fidelity and directional similarity. CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. CLIPD [^1] https://huggingface.co/docs/diffusers/conceptual/evaluation and View-CLIPD assess the consistency of transformations between two images, using text prompts similar to CLIP-Score and View-CLIP scores. Higher values are preferred. Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how HawkI++ stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches. See Supplementary Doc for additional results on $(-20^{\circ},210^{\circ}),(-20^{\circ},330^{\circ})\}$.
4. Limitations and Future Work
One limitation of our method is its reliance on an inference-time optimization routine for each scene and viewpoint, which may hinder real-time performance, achieving real-time capabilities for large, complex scenes and diverse, rich contents is a direction for future work. In addition, extending our approach to NVS and 3D applications with real-world constraints (correctness, safety requirements, etc), such as editing, object insertion, and composition, presents promising directions for further research.