### v2

# 3D-free meets 3D priors: Novel View Synthesis from a Single Image with Pretrained Diffusion Guidance

Abstract
Recent 3D novel view synthesis (NVS) methods are limited to single-object-centric scenes and struggle with complex environments. They often require extensive 3D data for training, lacking generalization beyond the training distribution. Conversely, 3D-free methods can generate text-controlled views of complex, in-the-wild scenes using a pretrained stable diffusion model without the need for a large amount of 3D-based training data, but lack camera control. In this paper, we introduce a method capable of generating camera-controlled viewpoints from a single input image, by combining the benefits of 3D-free and 3D-based approaches. Our method excels in handling complex and diverse scenes without extensive training or additional 3D and multiview data. It leverages widely available pretrained NVS models for weak guidance, integrating this knowledge into a 3D-free view synthesis approach to achieve the desired results. Experimental results demonstrate that our method outperforms existing models in both qualitative and quantitative evaluations, providing high-fidelity and consistent novel view synthesis at desired camera angles across a wide variety of scenes.
Figure 1: Our model can generate camera-controlled images at specified azimuth and elevation angles for diverse and complex scenes, without the need for additional 3D datasets or extensive training.
Figure 2: Method. Our method generates a high fidelity camera controlled novel viewpoint of a single image $I_{input}$, its text description and designated angle information. It infuses prior information from pre-trained NVS models into the text to image stable diffusion architecture in a 3D-free inference-time optimization procedure.
1 Introduction
Novel-view synthesis plays a pivotal role in numerous real-world applications, including 3D environments, augmented reality (AR), virtual reality (VR), and autonomous driving. Recent advancements in diffusion-based methods, such as Zero-1-to-3 (Zero123) Liu et al. (2023a) and Zero123++ Shi et al. (2023a), along with NeRFs Tancik et al. (2022); Zhou & Tulsiani (2023) and Gaussian Splatting Li et al. (2024); Zhu et al. (2023), have significantly propelled the field forward. These techniques enable the specification of camera angles and the sampling of novel-view images from precise viewpoints. However, diffusion-based methods remain largely object-centric and may struggle to generalize to complex scenes with intricate backgrounds. They also require extensive 3D object datasets for training. In contrast, NeRF and Gaussian Splatting methods can handle complex scenes but depend on multi-view information to construct 3D models. Therefore, achieving novel-view synthesis from a single image in a data-efficient manner, without relying on additional 3D, multi-view, or depth information, is highly advantageous.
Conversely, 3D-free methods like HawkI Kothandaraman et al. (2023) intelligently extract the rich 3D knowledge embedded in text-to-2D image diffusion models, such as stable diffusion Rombach et al. (2022); Podell et al. (2023), to generate text-controlled views from complex, real-world input images without needing additional multi-view or 3D information for fine-tuning or inference. However, these methods lack the ability to precisely control camera angles when generating novel-view images. Ideally, we aim for both data-efficient novel view synthesis and camera controllability, which is the primary focus of this paper.
We begin by analyzing how well CLIP understands the 3D space and the benefits of guidance images for 3D-free methods like HawkI. Using these insights, along with the established knowledge that 3D-based methods such as HawkI enable precise camera control and 3D-free methods like Zero123++ offer generalizability and data efficiency, we propose a simple approach for novel-view synthesis that generates camera-controlled novel views at specified azimuth and elevation angles from a single input image, without requiring 3D datasets or extensive training.
Our method utilizes information from off-the-shelf pretrained 3D-based models like Zero123++ Shi et al. (2023a) in conjunction with the pretrained stable diffusion model. It follows a 3D-free HawkI-style inference-time optimization procedure to achieve the desired results. We validate our approach through extensive qualitative and quantitative comparisons across various metrics that assess text consistency and fidelity.
To summarized our contributions, we (1) we analyze how well the CLIP model understands the 3D space and the usefulness of guidance image in 3D-free methods, (2) present a simple approach that uses 3D-priors within 3D-free 2D-based approaches for novel view synthesis, with precise camera control, that works well for complex image with multiple objects, backrgounds, etc.
2 Prior Work on 3D and 3D-free Approaches for NVS
Recent research has increasingly focused on novel view synthesis using diffusion models Chen et al. (2021); Mildenhall et al. (2021); Shi et al. (2021). Approaches for 3D generation Chen et al. (2024); Lin et al. (2023); Poole et al. (2022); Raj et al. (2023); Xu et al. (2023); Gao et al. (2024) often rely on text for reconstruction and require substantial multi-view and 3D data Shi et al. (2023b); Wang & Shi (2023); Yang et al. (2024); Liu et al. (2023b); Höllein et al. (2024) for supervised learning. For instance, Zero123 Liu et al. (2023a), Magic123 Qian et al. (2023), and Zero123++ Shi et al. (2023a) utilize a pre-trained stable diffusion model Rombach et al. (2022) combined with 3D data to learn various camera viewpoints. However, they are extremely data hungry, meaning that they need extensive multi-view and 3D data to train on. Additionally, most state-of-the-art view synthesis algorithms are largely object-centric, and may not work well on complex scenes containing multiple objects or background information. This is due to the domain gap between the 3D objects data that they are typically trained on, and the inferencing image.
Free3D Zheng & Vedaldi (2024) introduces an efficient method for synthesizing accurate 360-degree views from a single image without 3D representations. By incorporating the Ray Conditioning Normalization (RCN) layer into 2D image generators, it encodes the target view’s pose and enhances view consistency with lightweight multi-view attention layers and noise sharing. However, it still requires training on large-scale 3D datasets like Objaverse and multi-view information, and it cannot include background transformations. DreamFusion Poole et al. (2022) presents a Text-to-3D method using a NeRF and a Diffusion Model-based Text-to-2D model. It introduces a probability density distillation loss, allowing the 2D Diffusion Model to optimize image generation without needing 3D data or model modifications. DreamFusion’s key contributions are creating Text-to-3D models without 3D dataset training and utilizing a Diffusion Model. However, it cannot transform images with backgrounds or introduce elevation changes in camera-controlled images. HawkI Kothandaraman et al. (2023) synthesizes high-quality aerial view images using text and a single input image without 3D or multi-view information. It employs a pretrained text-to-2D Stable Diffusion model, achieving a balance between aerial view consistency and input image fidelity through test-time optimization and mutual information-based inference. However, HawkI struggles with controlling camera angles, detailed feature generation, and maintaining view consistency.
3 Understanding 2D models and the 3D space
3.1 How well do CLIP models understand the 3D-space?
This section examines how well CLIP models understand 3D space by focusing on the 3D-free method, HawkI Kothandaraman et al. (2023). HawkI uses classical computer vision techniques to generate aerial view images from ground view images via a homography transformation, which serves as the guidance image for the diffusion model. Our hypothesis is that HawkI’s ability to perform viewpoint transformations without 3D data relies on this homography image.
We conducted an experiment where the homography image was omitted to see if CLIP could generate camera-controlled images without it. Detailed angle instructions were still provided in the target text description. Our results showed that CLIP could not independently generate consistent viewpoints, highlighting the importance of 3D guidance images.
In the experiment, different pyramids, waterfalls, and houses were generated inconsistently, and camera control angles were not accurately followed. This demonstrates that CLIP struggles with 3D comprehension and validates the necessity of 3D guidance images. Figure 3 illustrates these findings.
Figure 3: Analysis of how well CLIP understands the 3D space In this experiment, we generate camera control images for specific angles without using any guidance image.
3.2 Importance of guidance image in 3D-free methods
Our previous analysis revealed that the CLIP model without a guidance image struggles to understand 3D space, resulting in inconsistent images. Conversely, HawkI with a guidance image cannot perform transformations from various camera viewpoints. This raises the question of what kind of guidance image is suitable for 3D-free camera control.
We conduct experiments using the images generated using the pretrained Zero123++ Shi et al. (2023a) model for guidance. In our experiments, the target text specified the desired transformation angles, but the guidance images had different angles. i.e. the guidance images introduced incorrect viewpoints. When generating an image at a $(30^{\circ},30^{\circ})$ angle, the model followed the guidance image’s suggestion, regardless of the text input. This indicates that the model benefits from the information in the 3D-derived guidance image, and that the guidance image is not simply a source of extra variance. Our experiment confirms the importance of guidance images in camera control, validating our methodology. Figure 4 illustrates these findings.
Figure 4: Using an image with an incorrect viewpoint as the guidance image In this experiment, we examine how the results are derived when an incorrect viewpoint image is used as a guidance image.
4 3D-free + 3D priors for Novel-View Synthesis
We present a simple method (Figure 2) for data-efficient text and camera-controlled novel-view synthesis from a single input image ($I_{input}$) and its text description ($t_{input}$) (e.g., “An ancient Egyptian pyramid in the desert,” obtained using the BLIP-2 Li et al. (2023) model). Our model eliminates the need for training data, 3D data, or multi-view data. Instead, it utilizes a pretrained text-to-2D image stable diffusion model as a strong prior, along with pretrained novel-view synthesis (NVS) models, e.g. Zero123++ for guidance. Our method combines information from the pretrained NVS model and performs a rapid inference-time optimization and inference routine, following a HawkI-style inference approach, to generate novel-view images of any given in-the-wild complex input scene at specified elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$).
Inference-time Optimization
We employ a pretrained NVS model $G$ to obtain a weak prediction, $I_{view}$, of $I_{input}$ at ($\alpha_{elev}$, $\alpha_{azi}$). This prediction is represented as $I_{view}=G(I_{input},\alpha_{elev},\alpha_{azi})$. Although $I_{view}$ is not a fully accurate depiction of the desired target, it provides weak or pseudo guidance for the model regarding the content and direction of the desired viewpoint transformation. Subsequently, we utilize the pretrained text-to-image stable diffusion model to perform inference-time optimization.
Specifically, initially, we enhance the CLIP embedding for $t_{input}$ with $I_{input}$ to derive $e_{optim}$ (optimized CLIP text-image embedding from $e_{input}$, which is the CLIP test embedding for $t_{input}$), which is optimized to most accurately reconstruct $I_{input}$:
[…]
Subsequently, the LoRA layers (with parameters $\theta_{LoRA}$) within the cross-attention layers of the diffusion model are fine-tuned at $e_{optim}$ to replicate $I_{input}$, employing the diffusion denoising loss function,
[…]
while keeping the rest of the UNet model unchanged. This process is then iterated for $I_{view}$, where $e_{optim}$ is again refined to $e_{view}$ to best reconstruct $I_{view}$. Following this, the LoRA layers are adjusted to capture the nuances of the weak guidance image $I_{view}$, guiding the transformation towards the desired outcome.
Thus, the predicted image from the 3D-based NVS method, Zero123++, serves as weak or pseudo guidance. The optimization strategy conditions the embedding space with knowledge related to the input image and its view variants. The guidance image prior increases the variance of the diffusion model, facilitating view transformation and providing a direction for viewpoint transformation.
Inference
To generate the camera-controlled image with designated elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$), we use the target text description $t_{target}$, which varies according to the corresponding $\alpha_{elev}$ and $\alpha_{azi}$. For instance, if $\alpha_{elev}=30^{\circ}$ and $\alpha_{azi}=30^{\circ}$, $t_{target}$ can be formatted as ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees” + $t_{input}$ (e.g., ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees, An ancient Egyptian pyramid in the desert.”). Next, we use the finetuned diffusion model to generate the target image using $t_{target}$, along with mutual information guidance Kothandaraman et al. (2023), which enforces similarity between the contents of the generated and input images.
5 Experiments and Results
Figure 5: Results on HawkI-Syn. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 6: Results on HawkI-Real. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 7: More results on HawkI-Syn. We present additional comparison results on HawkI-Syn for the angles of $(-20^{\circ},210^{\circ})$ and $(-20^{\circ},330^{\circ})$. Our model consistently produces view synthesis images that maintained background inclusion and view consistency, accurately mirroring the target elevation and azimuth angles. Notably, StableZero123 exhibits instability in its results. It’s important to highlight that this task specifically addresses negative azimuth angles. HawkI, for instance, fails to capture the correct camera elevation and is limited to generating aerial views. Zero123++ is capable of handling both elevation and azimuth but falls short in integrating background elements and intricate details, as also observed in previous outcomes. For example, when presented with an image of a pyramid casting a shadow, Zero123++ darkens the pyramid but fails to render the shadow accurately. This shortcoming is also apparent in images of a waterfall and a house. In the waterfall task within the specified azimuth range, Zero123++ produces an indistinct shape rather than a clear environment where water and lake are visible from below the rocks. Similarly, for the house image, it generates an incomplete image with gray patches. Conversely, our model not only captures the shadow details of the pyramid but also accurately renders the environment in the waterfall image, ensuring visibility of water and lake from beneath the rocks. Additionally, it adeptly incorporates details and backgrounds from multiple perspectives.
Figure 8: More Results on HawkI-Real. We extend our analysis to additional settings of $(-20^{\circ},210^{\circ})$ and $(-20^{\circ},330^{\circ})$. Our model, when tested on the HawkI-Real dataset, demonstrated superior performance in view synthesis images, excelling in background inclusion and view consistency, and accurately representing the target elevation and azimuth angles. In comparison to other leading models such as Zero123++, HawkI, and StableZero123, our model’s results are notably better. StableZero123’s outputs are incomplete, and Zero123++ struggles with capturing background details and intricate information. Specifically, Zero123++ neglected surrounding details, focusing solely on the Eiffel Tower. The original HawkI model also failed to achieve the correct camera elevation or produced images that overlooked important features. For example, in the cat transformation task, the output incorrectly depicted three cats instead of two. Our model stands out by delivering exceptional results for the Eiffel Tower, Hawaiian beach, and cat transformations, underscoring its advanced capabilities over other models. Furthermore, we present a quantitative evaluation in Table 2, which confirms our model’s dominance over state-of-the-art benchmarks across various metrics.
Datasets
We utilize the HawkI-Syn Kothandaraman et al. (2023) and HawkI-Real Kothandaraman et al. (2023) datasets that feature complex scenes with multiple foreground objects and background. Both datasets provide images and text prompts to the model. Explain the subset you used for your results, how you created the, how many images per sample you used for evaluation, etc.
Baselines
We compare our method with state-of-the-art view synthesis methods: Zero123++ Shi et al. (2023a) and Stable Zero123 for 3D-based methods and HawkI Kothandaraman et al. (2023) for 3D-free method.
Implementation Details
We employ the stable diffusion 2.1 model as the backbone for all our experiments and results. To generate the pseudo guidance images for different viewpoints, we use the pretrained Zero123++ Shi et al. (2023a) model. All images except those in HawkI-Real dataset are used at a resolution of $512\times 512$. For $I_{input}$, we train the text embedding for 1,000 iterations with the learning rate of $31e-$ 3 and the diffusion model for 500 iterations with the learning rate of $42e-$ 4. Training the text embedding for 1,000 iterations guarantees that the text embedding $e_{optim}$ is not too close to the $e_{input}$, avoiding bias towards $I_{input}$. Likewise, it is not too distant from $e_{input}$, allowing the text embedding space to capture the characteristics of $I_{input}$. Regarding $I_{view}$, we trained the text embedding for 500 iterations and the diffusion UNet for 250 iterations. We aim for $e_{view}$ to be near $e_{optim}$ and limit the diffusion model training to 250 iterations to prevent overfitting to $I_{view}$. The purpose of $I_{view}$ is to introduce variability and provide pseudo supervision rather than accurately approximating the camera control. We set the mutual information guidance hyperparameter to $61e-$ 6 and conduct inference over 50 steps.
5.1 Qualitative analysis
We evaluate our method on four distinct viewpoints:
$\{(\alpha_{elev},\alpha_{azi})\mid(30^{\circ},30^{\circ}),(-20^{\circ},210^{\circ}),(30^{\circ},270^{\circ}),\vspace*{-0.5em}(-20^{\circ},330^{\circ})\}.$
Our model is able to generate distinct viewpoints in the camera control to ensure consistency across generated views. Details are mentioned in the Zero123++ Shi et al. (2023a). We present qualitative representative results and comparisons with Zero123++ Shi et al. (2023a), HawkI Kothandaraman et al. (2023), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 5 and 6. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 5 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, our model accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints.
Similar observations are made for HawkI-Real results shown in Figure 6. StableZero123 is ineffective. Zero123++ fails to capture background or detailed information. For example, when tasked with camera control for an image of the Eiffel Tower, Zero123++ focuses solely on the Eiffel Tower, ignoring surrounding details. The original HawkI model, while producing aerial views, fails in angle conversion tasks. In contrast, our model accurately performs angle conversion tasks at $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$, including the Seine River in the background for the Eiffel Tower image, showcasing its superiority. Camera control tasks for the HawkI-Real dataset, including images like the Hawaii beach and a cat, further demonstrate our model’s excellence compared to other models. The key benefits of our model over 3D-based NVS methods such as Zero123++ and 3D-free methods such as HawkI arises by merging the strengths of 3D-based techniques into a 3D-free optimization process, effectively combining the best features of both.
5.2 Quantitative Evaluation
Following prior work Shi et al. (2023a); Kothandaraman et al. (2023); Liu et al. (2023a), we evaluate our method using six metrics - (i) LPIPS Zhang et al. (2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score Radford et al. (2021) and View-CLIP Score: Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{\text{elev}},\alpha_{\text{azi}})$ viewpoint, while the View-CLIP Score focuses specifically on the viewpoint. Higher values are preferred. (iii) CLIP-I Ruiz et al. (2023), CLIPD, and View-CLIPD: Evaluate image fidelity and directional similarity. CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. CLIPD [^1] https://huggingface.co/docs/diffusers/conceptual/evaluation and View-CLIPD assess the consistency of transformations between two images, using text prompts similar to CLIP-Score and View-CLIP scores. Higher values are preferred. Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 and Table 2 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how our models stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches.
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ View-CLIP $\uparrow$ CLIPD $\uparrow$ View-CLIPD $\uparrow$ CLIP-I $\uparrow$
HawkI-Syn $(30^{\circ},30^{\circ})$ Ours 0.5667 29.9125 19.7100 0.0534 0.0568 0.8334
HawkI-Syn $(30^{\circ},30^{\circ})$ HawkI 0.5998 27.6516 18.6122 0.0405 0.0549 0.8108
HawkI-Syn $(30^{\circ},30^{\circ})$ Zero123++ 0.5793 29.8709 18.4740 0.0552 0.0540 0.8330
HawkI-Syn $(30^{\circ},30^{\circ})$ Stable Zero123 0.6853 20.5335 18.0746 0.0528 0.0520 0.6181
HawkI-Syn $(30^{\circ},270^{\circ})$ Ours 0.5791 29.2654 19.1836 0.0553 0.0611 0.8274
HawkI-Syn $(30^{\circ},270^{\circ})$ HawkI 0.5996 29.0022 19.1088 0.0321 0.0498 0.8438
HawkI-Syn $(30^{\circ},270^{\circ})$ Zero123++ 0.5911 28.9917 19.1491 0.0482 0.0597 0.7917
HawkI-Syn $(30^{\circ},270^{\circ})$ Stable Zero123 0.6441 21.2231 18.3738 0.0406 0.0403 0.6152
HawkI-Real $(30^{\circ},30^{\circ})$ Ours 0.6075 30.9333 20.7728 0.0829 0.0730 0.8333
HawkI-Real $(30^{\circ},30^{\circ})$ HawkI 0.6405 28.5431 18.5430 0.0506 0.0468 0.8302
HawkI-Real $(30^{\circ},30^{\circ})$ Zero123++ 0.6679 29.8661 19.9530 0.0624 0.0467 0.8049
HawkI-Real $(30^{\circ},30^{\circ})$ Stable Zero123 0.6624 22.8604 18.7536 0.0523 0.0658 0.6740
HawkI-Real $(30^{\circ},270^{\circ})$ Ours 0.6180 30.0677 19.6756 0.0652 0.0709 0.8350
HawkI-Real $(30^{\circ},270^{\circ})$ HawkI 0.6497 28.4452 18.4231 0.0574 0.0577 0.8236
HawkI-Real $(30^{\circ},270^{\circ})$ Zero123++ 0.6935 27.2773 19.5926 0.0605 0.0698 0.7212
HawkI-Real $(30^{\circ},270^{\circ})$ Stable Zero123 0.6466 23.6915 17.8817 0.0593 0.0658 0.6726
Table 1: Quantitative Results. Evaluation on six metrics demonstrate the superior results of our method over prior work.
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ View-CLIP $\uparrow$ CLIPD $\uparrow$ View-CLIPD $\uparrow$ CLIP-I $\uparrow$
HawkI-Syn $(-20^{\circ},210^{\circ})$ Ours 0.5521 28.5599 17.9664 0.0502 0.0565 0.8535
HawkI-Syn $(-20^{\circ},210^{\circ})$ HawkI 0.5747 27.0915 16.9330 0.0403 0.0477 0.8370
HawkI-Syn $(-20^{\circ},210^{\circ})$ Zero123++ 0.7193 23.9836 16.7291 0.0496 0.0886 0.7389
HawkI-Syn $(-20^{\circ},210^{\circ})$ Stable Zero123 0.6938 20.3084 16.8253 0.0403 0.0400 0.5826
HawkI-Syn $(-20^{\circ},330^{\circ})$ Ours 0.5672 29.1569 18.1494 0.0556 0.0535 0.8656
HawkI-Syn $(-20^{\circ},330^{\circ})$ HawkI 0.5967 27.6743 17.3102 0.0520 0.0529 0.8409
HawkI-Syn $(-20^{\circ},330^{\circ})$ Zero123++ 0.6662 24.6859 17.5781 0.0688 0.0962 0.7383
HawkI-Syn $(-20^{\circ},330^{\circ})$ Stable Zero123 0.6201 22.3172 16.2088 0.0499 0.0542 0.7034
HawkI-Real $(-20^{\circ},210^{\circ})$ Ours 0.6222 29.6425 18.8914 0.0744 0.0709 0.7868
HawkI-Real $(-20^{\circ},210^{\circ})$ HawkI 0.6364 27.3844 17.5037 0.0568 0.0507 0.7821
HawkI-Real $(-20^{\circ},210^{\circ})$ Zero123++ 0.7739 24.3284 18.7046 0.0713 0.0951 0.6798
HawkI-Real $(-20^{\circ},210^{\circ})$ Stable Zero123 0.6637 21.2952 17.2609 0.0556 0.0704 0.5829
HawkI-Real $(-20^{\circ},330^{\circ})$ Ours 0.6135 30.6951 18.9849 0.0506 0.0371 0.8500
HawkI-Real $(-20^{\circ},330^{\circ})$ HawkI 0.6500 28.5102 17.4244 0.0481 0.0315 0.8463
HawkI-Real $(-20^{\circ},330^{\circ})$ Zero123++ 0.7132 25.5205 17.4653 0.0833 0.0977 0.7520
HawkI-Real $(-20^{\circ},330^{\circ})$ Stable Zero123 0.5730 23.9690 17.0343 0.0328 0.0726 0.7258
Table 2: Quantitative Results. Evaluation on six metrics demonstrate the superior results of our method over prior work.
6 Conclusions, Limitations and Future Work
In this paper, we propose an approach that integrates the advantages of off-the-shelf 3D-based pretrained models within 3D-free paradigms for novel view synthesis, offering precise control over camera angle and elevation. Our method performs effectively on complex, in-the-wild images containing multiple objects and background information. We qualitatively and quantitatively demonstrate the benefits of our method over corresponding 3D and 3D-free baselines. One limitation of our method is its reliance on an inference-time optimization routine for each scene and viewpoint, which may hinder real-time performance. Achieving faster performance is a direction for future work. Additionally, extending our approach to NVS and 3D applications with real-world constraints (such as respecting contact points and relative sizes) for tasks like editing, object insertion, and composition presents promising directions for further research. Based on the current results, we also propose exploring the use of an image-conditioned model to achieve a higher level of view consistency as a future research direction.
Appendix A Appendix
Figure 9: Additional comparisons in $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ settings on images from the HawkI-Syn and HawkI-Real datasets. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 10: Additional comparisons in $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ settings on images from the HawkI-Syn and HawkI-Real datasets. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 11: Additional comparisons in $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ settings on images from the HawkI-Syn and HawkI-Real datasets. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.
Figure 12: Additional comparisons in $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ settings on images from the HawkI-Syn and HawkI-Real datasets. Comparisons between the state-of-the-art view synthesis models, Zero123++, HawkI, Stable Zero123, and our method highlights the superior performance of our model in terms of background inclusion, view consistency, and the accurate representation of target elevation and azimuth angles.